Business and finance. When x number of units are sold, the price of each unit (in dollars) is given by p = -x/2 + 75. Find the unit price when the following quantities are sold: 2, 7, 9, 11.
Answer by sarah_adam(201)

Given that price of each unit is p = -x/2 + 75 where x is the number of units sold
we have to find price for the quantities 2,7,9 and 11
so in the equation of p = -x/2 +75 lets replace x with 2,7,9 and 11
price when 2 units are sold p = -(2)/2 + 75 ==> -1 + 75 = 74
therfore price for 2 units sold is $74
price when 7 units are sold p = -(7)/2 + 75 ==> -3.5 + 75 = 71.5 (Hint 7/2=3.5)
therfore price for 2 units sold is 71 dollars and 50 cents.
price when 9 units are sold p = -(9)/2 + 75 ==> -4.5 + 75 = 70.5
(Hint 9/2=4.5)
therfore price for 2 units sold is 70 dollars and 50 cents

price when 11 units are sold p = -(11)/2 + 75 ==> -5.5+ 75 = 69.5
(Hint 11/2=5.5)
therfore price for 11 units sold is 69 dollars and 50 cents